Price = 12,15,18,25,22,18,30.
Demand = 65,60,50,41,40,56,45.
Required following.
1) coefficient of correlation
2) coefficient of determination
3) probable Error
4) Interpretation
1
Expert's answer
2020-05-07T15:52:10-0400

1) X Values: ∑ = 140,Mean = 20

∑(X - Mx)2 = SSx = 226

Y Values: ∑ = 357,Mean = 51

∑(Y - My)2 = SSy = 560

X and Y Combined

N = 7

∑(X - Mx)(Y - My) = -297

r = ∑((X - My)(Y - Mx)) / "\sqrt{((SSx)(SSy))}"

r = -297 / "\sqrt{((226)(560))}" = -0.8348

2) cd=r2=0.69697

3)using formula (where 0.674 - const): e=0.674 *(1-r2) /"\sqrt{N}" =0.0771

​4)A strong downhill (negative) linear relationship